# FOSS: A Self-Learned Doctor for Query Optimizer

### v1

@@ Abstract
Various works have utilized deep reinforcement learning (DRL) to address the query optimization problem in database system. They either learn to construct plans from scratch in a bottom-up manner or guide the plan generation behavior of traditional optimizer using hints. While these methods have achieved some success, they face challenges in either low training efficiency or limited plan search space. To address these challenges, we introduce FOSS, a novel DRL-based framework for query optimization. FOSS initiates optimization from the original plan generated by a traditional optimizer and incrementally refines suboptimal nodes of the plan through a sequence of actions. Additionally, we devise an asymmetric advantage model to evaluate the advantage between two plans. We integrate it with a traditional optimizer to form a simulated environment. Leveraging this simulated environment, FOSS can bootstrap itself to rapidly generate a large amount of high-quality simulated experiences. FOSS then learns and improves its optimization capability from these simulated experiences. We evaluate the performance of FOSS on Join Order Benchmark, TPC-DS, and Stack Overflow. The experimental results demonstrate that FOSS outperforms the state-of-the-art methods in terms of latency performance and optimization time. Compared to PostgreSQL, FOSS achieves savings ranging from 15% to 83% in total latency across different benchmarks.
@@
query optimization, deep reinforcement learning, DBMS
@@ I Introduction
However, for traditional query optimizer, finding the optimal plan within the vast plan space is an NP-hard problem [2]. Even if only the left-deep tree is considered in join order, the plan space still grows exponentially to $O(n!)$, where $n$ represents the number of tables involved in the query. The plan search space is further expanded when considering other factors, such as bushy tree, join methods and access path. Meanwhile, errors in traditional cardinality and cost estimator often result in poor-performance plans [2]. Furthermore, traditional query optimizer can not learn from the past experiences, often generating the similar poor-performance plans for the same queries repeatedly [3].
@@ I-A Machine Learning for Query Optimization
To address the problem of traditional query optimizer, numerous studies focus on employing machine learning in end-to-end query optimization. We roughly categorize current methods into two types: plan-constructor and plan-steerer.
Plan-constructor. It is an expertise-omission approach that discards or underuses expert knowledge of traditional optimizer and focuses on constructing plans from scratch [4, 5, 6, 7, 8, 9]. It models the process of query plan generation as a Markov Decision Process (MDP) and uses deep reinforcement learning (DRL) to solve it. Its framework involves several components. An agent (i.e., a learned model) starts with a set of individual tables involved in the query, and at each step, it joins tables, joins subplans or specifies an operator. Each step results in a new state, which consists of a partial plan and other tables that have not been joined yet. The process continues until all tables involved in the query are joined, resulting in a complete plan. In this process, the DBMS serves as the environment and provides reward feedback, which are typically correlated with the execution cost or true execution latency of the final plans. Although plan-constructor can gradually surpass the performance of traditional query optimizer after a certain training period, it has the following drawbacks:
@@
  • (C3) Sparse reward. The number of steps required for an agent to complete the construction of a plan depends on the number of tables involved in the query. In Join Order Benchmark (JOB) [2], the queries have between 3 and 16 joins, with an average of 8 joins per query. However, prior methods such as [8, 7, 6, 10] encounter challenges in providing specific evaluations for subplans before obtaining the complete plan. In their scenarios, rewards are often assigned a value of 0 for the intermediate state, leading to sparse rewards for the agent and limiting access to sufficient guidance [11].
Plan-steerer. It is a blackbox-expertise approach that leverages the embedded expert knowledge in traditional query optimizer [12, 13, 14]. Guided by various hints, it generates multiple plans through the traditional query optimizer. For instance, in Bao [12], coarse-grained hints like disabling nested loop join for the entire query are used to guide the traditional query optimizer. It then utilizes a learned value network to evaluate these candidate plans and predict the cost. Ultimately, it selects the plan with the potentially lowest cost for execution. However, it has the following problems:
@@ I-B A Novel Framework for Query Optimization
However, to obtain an excellent plan-doctor, two key capabilities are essential: i) the ability to identify suboptimal aspects in the original plan and generate appropriate candidate action sequences for optimization. Note that different action sequences applied to the original plan will result in distinct new plans. ii) the ability to select the sequence that yields the optimal result for the original plan from candidates. To achieve these objectives, for i), FOSS learns a planner that models the identification of suboptimal aspects and the generation of action sequences as a MDP and employs DRL to solve it, and for ii), FOSS learns an asymmetric advantage model (AAM) to serve as the optimal action sequence selector. AAM determines the optimal sequence by evaluating the advantage values between pairs of candidate plans generated through candidate action sequences. Furthermore, inspired by model-based reinforcement learning (MBRL), FOSS integrates the AAM with a traditional optimizer to form a simulated environment, where the AAM serves as a reward indicator, and the traditional optimizer acts as a state transitioner.
@@ I-C Contributions
  • We design an asymmetric advantage model to serve as a reward indicator in the simulated environment and as a selector to choose the optimal plan.
  • Experimental results show that FOSS outperforms the state-of-the-art (SOTA) methods in terms of latency performance and optimization time, while also reducing total latency by 15% to 83% compared to the PostgreSQL.
@@ II SYSTEM OVERVIEW
Planner. We deploy the entire process of planner using DRL framework. Starting from the original plan, planner first encodes plan features. Considering the temporal nature of plans generated by planner, we incorporate the step status. These features are concatenated as the initial state and then passed to the agent. The agent first represents the state using a state network. Then the state representation is subsequently forwarded to the action selector to generate an action, which could involve operations such as swapping table positions or overriding join method. Following the execution of the action, a new plan is generated and forwarded to the reward function module to obtain rewards. Meanwhile, the new plan is utilized to generate a new state. This iterative process continues until the maximum step limit is attained. Experiences in the format $\{State,Action,Reward,State_{{}^{\prime}}\}$ will be collected to update the agent. Through extensive training, planner can generate more optimal plans than the original plan.
@@
Asymmetric Advantage Model. It serves as both optimal plan selector and the reward indicator in simulated environment. It consists of state network and position-aware output layer. First, we define the state network, which also helps the planner’s agent understand the state. Its primary architecture is based on the Transformer [15], which can efficiently capture the structural and node information of plans. It involves embedding plan features and step statuses separately, passing them through a multi-head attention network, ultimately outputting the state representations. For position-aware output layer, the input comprises the two state representations $statevec_{left}$ and $statevec_{right}$, while the output indicates the extent to which the $statevec_{right}$ surpasses the $statevec_{left}$. We map the output to different scores. The selection of the optimal plan and the evaluation of rewards is determined based on these scores.
@@
Simulated Learner. In DBMS, the cost of using the traditional optimizer to generate plans is low. FOSS utilizes the traditional optimizer and AAM to form a simulated environment. Therefore, FOSS comprises three key components: the agent, the real environment, and the simulated environment. The agent of planner interacts with the real environment, which executes explored plans in the actual system to gather AAM training pool. These data are then used to update the AAM through a supervised learning process. Then, the agent will interact with the simulated environment to obtain simulated experiences, which will be used to update the agent. After several simulated learning iterations, we use the updated agent to continue interacting with the real environment, repeating the above process. Furthermore, we incorporate a dynamics timeout mechanism to ensure efficient data collection during the interaction with the real environment. We also integrate a validation mechanism. During simulated learning iterations, we select plans with high estimated performance, execute these plans to measure their execution latency, and then incorporate them into the AAM training pool.
@@ III Planner
Given a schema with $n$ tables, when a query $Q$ is received, we obtain the complete plan $CP$ (i.e., execution plan) from the output of the traditional optimizer. We extract the key information from $CP$ that holds significant importance on the performance of plan. This yields an incomplete plan $ICP$, as shown in Fig. 2. The $ICP$ represents a subset of $CP$ and includes the join order and join method. The planner’s main process is to iteratively modify $ICP$ in a step-by-step manner. After taking one step on the $ICP$, guided by the new $ICP$, we can derive the new $CP$ using the DBMS hook, such as pg_hint_plan in PostgreSQL. When the maximum number of steps $maxsteps$ per episode is reached, the plan modification process will be stopped. Planner adopts the framework of DRL, which consists of key components like agent, environment, state, action and reward. Next, we will provide a detailed explanation of each component.
State. It is the output of the environment and the input to the agent. It is primarily composed of $CP$, typically in the form of tree-like structure. To facilitate the agent’s handling of $CP$, it is essential to extract critical information from the plan and encode it. We will provide the specific details of plan encoding in IV-A. Now, we use $PlanEncoding(CP)$ to represent the encoding of $CP$. Additionally, to incorporate the temporal characteristics of the state, we also include the $Step(t)=t/maxsteps$ in state. By concatenating these two types of features, we obtain the state features $State(CP_{t})$ at time step $t$.
@@
Next, we will explain the representation of two types of action. The first type is swapping the positions of two tables $(T_{l},T_{r})$, represented as $Swap(T_{l},T_{r})$. There are $num_{s}=(n\times(n-1))/2$ different actions for table positions swapping. The second type is overriding the join method of node $O_{i}$ with the $j$-th join method in $Op$, represented as $Override(O_{i},Op_{j})$, where $Op$ is the set of all available join methods in DBMS. There are $num_{o}=\left|Op\right|\times(n-1)$ different actions for overriding join method. Action is encoded as an integer $a$ within the range $[1,num_{s}+num_{o}]$. We define an action mapping function $Act(a,ICP)$ which maps an integer $a$ to the corresponding action behavior on $ICP$. We can represent it as follows:
$Act(a,ICP)=\begin{cases}Swap(T_{l},T_{r})&a\in Range_{s}\\ Override(O_{i},Op_{j})&a\in Range_{o}\end{cases}$ (2)
, where the integer in $Range_{s}=[1,num_{s}]$ represents table positions swapping action, and the integer in $Range_{o}=[num_{s}+1,num_{s}+num_{o}]$ represents join method overriding action. If $a$ satisfies $Range_{s}$, the $Swap(T_{l},T_{r})$ is called, where $l$ and $r$ satisfy the following conditions:
@@
. If $a$ satisfies $Range_{o}$, the $Override(O_{i},Op_{j})$ is called, where $i$ and $j$ satisfy the following conditions:
$\begin{cases}i=\lceil\frac{num_{s}+num_{o}+1-a}{\left|Op\right|}\rceil\\ j=(num_{s}+num_{o}-a\mod\left|Op\right|)+1\end{cases}$
@@
. However, each query may have different action space restrictions, depending on the number of tables in the query graph. For example, in the query shown in Fig. 2, $Swap(T_{1},T_{5})$ is considered an illegal action. To address the problem, we perform a validity check on the action space before the agent selects an action. We use $actionmask$ to prohibit illegal actions. To further prune the action space, we employ a heuristic rule. After execute $Swap(T_{l},T_{r})$, the subsequent action can only select how to override the join method related to the two tables that were swapped in the previous step. This also be implemented using the $actionmask$.
Reward. During the training process of planner, we maintain both the known best plan $CP^{q}_{b}$ for each query and the best plan $CP^{p}_{b}$ perceived by FOSS. And at each step $t$ of one episode $e$, a new incomplete plan $ICP_{t}^{e}$ and a new complete plan $CP_{t}^{e}$ will be generated. We determine reward based on these plans. The reward at each step $Reward_{t}^{e}$ is the sum of bounty $Bounty_{t}^{e}$ and penalty $Penalty_{t}^{e}$. And the reward for each episode is the sum of reward obtained at each step.
@@
Bounty. First, we define the initial advantage function $Adv_{initial}$ to calculate the initial advantage value between two plans, indicating how much better $CP_{r}$ is compared to $CP_{l}$.
$Adv_{initial}(CP_{l},CP_{r})=\frac{U(CP_{l})-U(CP_{r})}{U(CP_{l})}$
@@
, where $U(CP)$ is related to the performance of $CP$. However, fine-grained value is unsuitable for dynamically changing system workloads. We discretize initial advantage value into several intervals. We take an $l$-element finite ordered point set $\{d_{i}|i\in\{1,2,...,l\},0\leq d_{1}<...<d_{l}<1\}$ and use it to partition the interval $(-\infty,1]$ into $l+1$ subintervals, resulting in the set of ordered subintervals $D=\{(d_{i},d_{i+1}]|i\in\{0,1,...,l\},d_{0}\to-\infty,d_{l+1}=1\}\}$. We use $D_{k}$ to represent the $k$-th element (i.e., subinterval) in $D$ and set $\hat{D}_{k}=(d_{k}+d_{k-1})/2$. We define the final advantage function $Adv$ that takes a pair of plans as input and outputs a score.
$Adv(CP_{l},CP_{r})=k-1\ \text{ if }Adv_{initial}(CP_{l},CP_{r})\in D_{k}$ (3)
In each episode, we maintain the current best plan $CP_{b}^{e}$. At each step, we calculate $Adv(CP_{b}^{e},$ $CP_{t}^{e})$ and provide a step bounty $cb_{t}^{e}$:
$cb_{t}^{e}=Adv(CP_{b}^{e},CP_{t}^{e})\times\eta,\text{ $\eta>0$}$
. Additionally, if current step is $maxsteps$, we will append the episode bounty $eb^{e}$. We begin by defining $eb^{pb}$ and $eb^{qb}$ to represent the FOSS’s episode bounty and known best episode bounty.
$eb^{pb}=Adv_{initial}(CP^{q}_{o},CP^{p}_{b})\times eb^{m}$
$eb^{qb}=Adv_{initial}(CP^{q}_{o},CP^{q}_{b})\times eb^{m}$
, where $eb^{m}$ represents the maximum episode bounty. Then we calculate the advantage values $adv_{1}=Adv(CP^{q}_{b},CP_{b}^{e})$, $adv_{2}=Adv(CP^{p}_{b},CP_{b}^{e})$, and $adv_{3}=Adv(CP^{q}_{o},CP_{b}^{e})$. The episode bounty $eb^{e}$ can be represented as follows:
$eb^{e}=\begin{cases}\hat{D}_{adv_{1}+1}\times(eb^{m}-eb^{qb})+eb^{qb}&adv_{1}>0\\ \hat{D}_{adv_{2}+1}\times(eb^{qb}-eb^{pb})+eb^{pb}&adv_{1}=0,adv_{2}>0\\ \frac{adv_{3}}{l}\times eb^{pb}&adv_{1},adv_{2}=0\end{cases}$
. Finally, we obtain the $Bounty_{t}^{e}$ as follows:
$Bounty_{t}^{e}=cb_{t}^{e}+\lfloor\frac{t}{maxsteps}\rfloor\times eb^{e}.$
@@
Penalty. To encourage the agent to reach each different state with as few steps as possible, we implement a penalty measure. In planner, reaching a certain state from the initial state can involve multiple action sequences of varying lengths, where an action sequence refers to a combination of $Swap$ and $Override$ operators. In order to gain more rewards, the agent may opt for inappropriate action sequences. For example, assume that the execution latency for the hash join operator at position $O_{i}$ is $L_{1}$, $L_{2}$ for merge join, and $L_{3}$ for nested loop join. Suppose $L_{1}>L_{2}>L_{3}$, and the original plan choose using hash join. In order to obtain more bounty, the agent may choose to initially override the method at $O_{i}$ to a merge join, and then in subsequent steps, override it to a nested loop join. To address such issue, we calculate the minimum number of steps required from the original incomplete plan to the current incomplete plan, denoted as $minsteps(ICP_{t}^{e})$. Additionally, we introduce a penalty coefficient $\gamma>0$. We set the penalty value according to the following formula:
@@
. In the case of the initial step, it takes the query $Q$ as input and outputs a complete plan $CP$. In non-initial steps, it takes both the incomplete plan $ICP$ and the query $Q$ as input, generating a complete plan $CP$ steered by $ICP$. We can then encode the $CP$ to obtain the new state by using (1). For ii), after obtaining $CP$, we provide a reward as described in the part of Reward. The key difference is that in the real environment, we use the DBMS executor $\Psi_{p}$ to execute the plan and then use the execution latency to represent the performance of plan $U(CP)$. Subsequently, we calculate the advantage and the final reward. In the simulated environment, the advantage value is evaluated by asymmetric advantage model without executing the plan. The specific details will be discussed in IV-B.
@@
2:     $CP_{0}\leftarrow\Gamma_{p}(Q,/)$
3:     $ICP_{0}\leftarrow Extract(CP_{0})$
4:     HistoryPlanSet $T=[ICP_{0}]$
5:     EstimatedBestPlan $CP_{best}=CP_{0}$
@@
7:         $actionmask\leftarrow ICP_{t-1}$
8:         if $t>1\cap a_{t-1}\in Range_{s}$ then
@@
11:         $s_{t-1}=State(CP_{t-1})$
12:         $statevec_{t-1}=\phi(s_{t-1})$
13:         $a_{t}=\pi(statevec_{t-1},actionmask)$
14:         $ICP_{t}=Act(a_{t},ICP_{t-1})$
15:         $CP_{t}\leftarrow\Gamma_{p}(Q,ICP_{t})$
16:         $r_{t}=Penalty_{t}$
17:         if $ICP_{t}\notin T$ then
18:              $r_{t}=r_{t}+Bounty_{t}$
19:              $Add\ ICP_{t}\ to\ T$
@@
21:         $CP_{best}\leftarrow CP_{t}\text{\ \ if\ \ }Adv(CP_{best},CP_{t})>0$
22:         $experiences\leftarrow Collect(s_{t-1},a_{t},r_{t},s_{t})$

### v2

@@ Abstract
Various works have utilized deep learning to address the query optimization problem in database system. They either learn to construct plans from scratch in a bottom-up manner or steer the plan generation behavior of traditional optimizer using hints. While these methods have achieved some success, they face challenges in either low training efficiency or limited plan search space. To address these challenges, we introduce FOSS, a novel framework for query optimization based on deep reinforcement learning. FOSS initiates optimization from the original plan generated by a traditional optimizer and incrementally refines suboptimal nodes of the plan through a sequence of actions. Additionally, we devise an asymmetric advantage model to evaluate the advantage between two plans. We integrate it with a traditional optimizer to form a simulated environment. Leveraging this simulated environment, FOSS can bootstrap itself to rapidly generate a large amount of high-quality simulated experiences. FOSS then learns from these experiences to improve its optimization capability. We evaluate the performance of FOSS on Join Order Benchmark, TPC-DS, and Stack Overflow. The experimental results demonstrate that FOSS outperforms the state-of-the-art methods in terms of latency performance. Compared to PostgreSQL, FOSS achieves speedup ranging from 1.15x to 8.33x in total latency across different benchmarks.
@@
query optimization, deep learning, DBMS
@@ I Introduction
However, for traditional query optimizer, finding the optimal plan within the vast plan space is an NP-hard problem [2]. Even if only the left-deep tree is considered in join order, the plan space still grows exponentially to $O(n!)$, where $n$ represents the number of tables involved in the query. The plan search space is further expanded when considering other factors, such as bushy tree, join methods and access path. Meanwhile, errors in traditional cardinality and cost estimator often result in poor-performance plans [2, 3]. Furthermore, traditional query optimizer can not learn from the past experiences, often generating the similar poor-performance plans for the same queries repeatedly [4].
@@ I-A Machine Learning for Query Optimization
To address the problem of traditional query optimizer, numerous studies focus on employing deep learning in end-to-end query optimization. We roughly categorize current methods into two types: plan-constructor and plan-steerer.
Plan-constructor. It is an expertise-omission approach that discards or underuses expert knowledge of traditional optimizer and focuses on constructing plans from scratch [5, 6, 7, 8, 9, 10]. It models the process of query plan generation as a Markov Decision Process (MDP) and uses deep reinforcement learning (DRL) to solve it. Its framework involves several components. An agent (i.e., a learned model) starts with a set of individual tables involved in the query, and at each step, it joins tables, joins subplans or specifies an operator. Each step results in a new state, which consists of the partial plan and other tables that have not been joined yet. The process continues until all tables involved in the query are joined, resulting in a complete plan. In this process, the DBMS serves as the environment and provides reward feedback, which are typically correlated with the execution cost or true execution latency of the final plans. Although plan-constructor can gradually surpass the performance of traditional query optimizer after a certain training period, it has the following drawbacks:
@@
  • (C3) Sparse reward. The number of steps required for an agent to complete the construction of a plan depends on the number of tables involved in the query. In Join Order Benchmark (JOB) [2], the queries have between 3 and 16 joins, with an average of 8 joins per query. However, prior methods such as [8, 7, 11] encounter challenges in providing specific evaluations for subplans before obtaining the complete plan. In their scenarios, rewards are often assigned a value of 0 for the intermediate state, leading to sparse rewards for the agent and limiting access to sufficient guidance [12].
Plan-steerer. It is a blackbox-expertise approach that leverages the embedded expert knowledge in traditional query optimizer [13, 14, 15, 16]. Guided by various hints, it generates multiple plans through the traditional query optimizer. For instance, in Bao [13], coarse-grained hints like disabling nested loop join for the entire query are used to guide the traditional query optimizer. It then utilizes a learned value network to evaluate these candidate plans and predict the cost. Ultimately, it selects the plan with the potentially lowest cost for execution. However, it has the following problems:
@@ I-B A Novel Framework for Query Optimization
However, to obtain an excellent plan-doctor, two key capabilities are essential: i) the ability to identify suboptimal aspects in the original plan and generate appropriate candidate action sequences for optimization. Note that different action sequences applied to the original plan will result in distinct new plans. ii) the ability to select the sequence that yields the optimal result for the original plan from candidates. To achieve these objectives, for i), FOSS learns a planner that models the identification of suboptimal aspects and the generation of action sequences as a MDP and employs DRL to solve it, and for ii), FOSS learns an asymmetric advantage model (AAM) to serve as the optimal candidate action sequence selector. AAM evaluates advantage values between pairs of candidate plans generated by candidate action sequences to determine the optimal plan among them. Furthermore, inspired by model-based reinforcement learning (MBRL), FOSS integrates the AAM with a traditional optimizer to form a simulated environment, where the AAM serves as a reward indicator, and the traditional optimizer acts as a state transitioner.
@@ I-C Contributions
  • We design an asymmetric advantage model to serve as a reward indicator in the simulated environment and as a selector to choose the optimal plan among candidate plans.
  • Experimental results show that FOSS outperforms the state-of-the-art methods in terms of latency performance, while also achieves speedup ranging from 1.15×1.15$\times$1.15 × to 8.33×8.$3$ in total latency compared to the PostgreSQL.
@@ II SYSTEM OVERVIEW
Planner. We deploy the entire process of planner using DRL framework. Starting from the original plan, planner first encodes plan features. Considering the temporal nature of plans generated by planner, we incorporate the step status. These features are concatenated as the initial state and then passed to the agent. The agent first represents the state using a state network. Then the state representation is subsequently forwarded to the action selector to generate an action, which could involve operations such as swapping table positions or overriding join method. Following the execution of the action, a new plan is generated and forwarded to the reward function module to obtain rewards. Meanwhile, the new plan is utilized to generate a new state. This iterative process continues until the maximum step limit is attained. Experiences in the format $\{State,Action,Reward,State_{{}^{\prime}}\}$ will be collected to update the agent. Through extensive training, planner can generate better plans than the original plan.
@@
Asymmetric Advantage Model. It serves as both optimal plan selector and reward indicator in simulated environment. It consists of state network and position-aware output layer. First, we define the state network, which also helps the planner’s agent analyze the state. Its primary architecture is based on the Transformer [17], which can efficiently capture the structural and node information of plans. It involves embedding plan features and step statuses separately, passing them through a multi-head attention network, ultimately outputting the state representations. For position-aware output layer, the input comprises two state representations $statevec_{left}$ and $statevec_{right}$, while the output indicates the extent to which $statevec_{right}$ surpasses $statevec_{left}$. We map the output to different scores. The selection of the optimal plan and the evaluation of rewards is determined based on these scores.
@@
Simulated Learner. In DBMS, the cost of using the traditional optimizer to generate plans is low. FOSS utilizes the traditional optimizer and AAM to form a simulated environment. Therefore, FOSS comprises three key components: the agent, the real environment, and the simulated environment. During the training of FOSS, the agent will concurrently interact with both the real environment and the simulated environment. During the interaction with the real environment, the generated plans are executed in DBMS to obtain execution latency, forming the AAM training data pool. The data are then used to update the AAM through a supervised learning process. We also incorporate a dynamics timeout mechanism to ensure efficient execution latency collection. During the interaction with the simulated environment, a large amount of high-quality simulated experiences will be generated to update the agent. Meanwhile, we integrate a validation mechanism during the collection of simulated experiences. FOSS will collect the plans with high estimated performance, execute these plans in real environment to obtain their execution latency, and then incorporate them into the AAM training pool.
@@ III Planner
Given a schema with $n$ tables, when a query $q$ is received, we obtain the complete plan $CP$ (i.e., execution plan) from the output of the traditional optimizer. As shown in the top right corner of Fig. 2, we extract the join methods and join order from $CP$, which significantly affect the performance of $CP$ [10]. We refer to such a tree structure containing only the join order and join methods as the incomplete plan $ICP$. The planner’s main process is to iteratively modify $ICP$ in a step-by-step manner. After taking one step on the $ICP$, guided by the new $ICP$, we can derive the new $CP$ using the DBMS extension, such as pg_hint_plan[18] in PostgreSQL. With the extension, table scan operators and other nodes will be complemented by the traditional optimizer using its own expert knowledge, resulting in the execution plan $CP$. When the maximum number of steps $maxsteps$ is reached, the plan modification process will be stopped. We refer to this process as an episode for an input query. Planner adopts the framework of DRL, which consists of key components like agent, environment, state, action and reward. Next, we will provide a detailed explanation of each component.
State. In DRL, it is the output of the environment and the input to the agent. In planner, it is primarily composed of $CP$. To facilitate the subsequent processing by the agent, we pick up the key features from $CP$ and encode them into vectors. We will provide the specific details of $CP$ encoding in IV-A. Now, we use $PlanEncoding(CP)$ to represent the encoding of $CP$. Additionally, to incorporate the temporal characteristics of the state, we also include the $Step(t)=t/maxsteps$ in state. By concatenating these two types of features, we obtain the state features $State(CP_{t})$ at time step $t$.
@@
Next, we will explain the representation of two types of action. The first type is swapping the positions of two tables $(T_{l},T_{r})$, represented as $Swap(T_{l},T_{r})$. There are $I_{s}=(n\times(n-1))/2$ different actions for table positions swapping. The second type is overriding the join method of node $O_{i}$ with the $j$-th join method in $Op$, represented as $Override(O_{i},Op_{j})$, where $Op$ is the set of all available join methods in DBMS. There are $I_{o}=\left|Op\right|\times(n-1)$ different actions for overriding join method. Action is encoded as an integer $a$ within the range $[1,I_{s}+I_{o}]$. We define an action mapping function $Act(a,ICP)$ which maps an integer $a$ to the corresponding action behavior on $ICP$. We can represent it as follows:
$Act(a,ICP)=\begin{cases}Swap(T_{l},T_{r})&a\in[1,I_{s}]\\ Override(O_{i},Op_{j})&a\in[I_{s}+1,I_{s}+I_{o}]\end{cases}$
. When $Swap(T_{l},T_{r})$ is called, $l$ and $r$ satisfy the following conditions:
@@
.When $Override(O_{i},Op_{j})$ is called, $i$ and $j$ satisfy the following conditions:
$\begin{matrix}i=\lceil\frac{I_{s}+I_{o}+1-a}{\left|Op\right|}\rceil\\ j=(I_{s}+I_{o}-a\mod\left|Op\right|)+1\end{matrix}$
@@
. However, each query may have different action space restrictions, depending on the number of tables in the query graph. For example, in the query shown in Fig. 2, $Swap(T_{1},T_{5})$ is considered an illegal action. To address the problem, we perform a validity check on the action space before the agent selects an action. We use $actionmask$ to prohibit illegal actions. To further prune the action space, we employ a heuristic rule. After executing $Swap(T_{l},T_{r})$, the subsequent action can only be $Override(O_{i},Op_{j})$, where $O_{i}$ is the parent node of $T_{l}$ or $T_{r}$. This also be implemented using the $actionmask$.
Reward. The design of the reward function plays a crucial role in DRL. It evaluates the effect of actions chosen by the agent. In planner, the reward $Reward_{t}^{e}$ at each step $t$ of one episode $e$ is composed of the sum of $Bounty_{t}^{e}$ and $Penalty_{t}^{e}$. The former provides positive feedback for actions selected by the agent, while the latter penalizes inappropriate actions.
@@
Bounty. First, we define the initial advantage function $Adv_{init}$ to calculate the initial advantage between two plans, indicating how much better $CP_{r}$ is compared to $CP_{l}$.
$Adv_{init}(CP_{l},CP_{r})=\frac{U(CP_{l})-U(CP_{r})}{U(CP_{l})}\in(-\infty,1)$
@@
, where $U(CP)$ is related to the performance of $CP$ (e.g., latency). However, fine-grained value is unsuitable for dynamically changing system workloads. We discretize initial advantage into several intervals. We take an $l$-element finite ordered point set $\{d_{i}|i\in\{1,2,...,l\},0\leq d_{1}<...<d_{l}<1\}$ and use it to partition the interval $(-\infty,1]$ into $l+1$ subintervals, resulting in the set of ordered subintervals $D=\{(d_{i},d_{i+1}]|i\in\{0,1,...,l\},d_{0}\to-\infty,d_{l+1}=1\}$. We use $D_{k}$ to represent the $k$-th element in $D$ and set $\hat{D}_{k}=(d_{k}+d_{k-1})/2$, $\hat{D}_{0}=0$. We define the final advantage function $Adv$ that takes a pair of plans as input and outputs a score.
$Adv(CP_{l},CP_{r})=k-1\ \text{ if }Adv_{init}(CP_{l},CP_{r})\in D_{k}$ (2)
For each episode $e$ of query $Q$, the planner starts from the original plan $CP^{Q}_{ORI}$ and aims to generate a better plan $CP_{t}^{e}$ than the plans from previous $t-1$ steps at step $t$. We assign the step-bounty $pb_{t}^{e}=Adv(\overline{CP}^{e}_{t-1},CP_{t}^{e})$, where $\overline{CP}^{e}_{t-1}$ is the estimated optimal plan among all plans from the previous $t-1$ steps. Naturally, when an episode $e$ concludes, the final output plan $\overline{CP}^{e}$ generated by FOSS (i.e., the plan to be executed) should carry significant weight in the reward. Therefore, we add the episode-bounty $eb^{e}$ based to the performance of $\overline{CP}^{e}$. To comprehensively evaluate the performance of $\overline{CP}^{e}$, we select the best-performing plan and the median-performing plan from the set of executed plans for query Q that outperform the original plan. These plans, along with the original plan, form the reference plan set $CP^{ref}$. Then we calculate the actual bounty of each reference plan using the formula $refb_{i}=Adv_{init}(CP^{Q}_{ORI},CP^{ref}_{i})$ to serve as reference bounty. Finally, we estimate the $adv_{i}=Adv(CP^{ref}_{i},\overline{CP}^{e})$ to determine the ranking range of $\overline{CP}^{e}$. The $eb^{e}$ is defined as follows:
$eb^{e}=\sum_{i=1}^{3}(\hat{D}_{adv_{i}}+\frac{adv_{i}}{l})\times(refb_{i-1}-refb_{i})$
, where $refb_{0}$ is set to 1, indicating the upper limit value. $Bounty_{t}^{e}$ is represented as follows:
$Bounty_{t}^{e}=pb_{t}^{e}+\eta\times\lfloor\frac{t}{maxsteps}\rfloor\times eb^{e}$
, where $\eta$ is a constant representing the weight of episode-bounty relative to step-bounty.
@@
Penalty. To encourage the agent to reach each different state with as few steps as possible, we implement a penalty measure. In planner, reaching a certain state from the initial state can involve multiple action sequences of varying lengths, where an action sequence refers to a combination of $Swap$ and $Override$ operators. In order to gain more rewards, the agent may opt for inappropriate action sequences. For example, assume that the execution latency for the hash join operator at position $O_{i}$ is $L_{1}$, $L_{2}$ for merge join, and $L_{3}$ for nested loop join. Suppose $L_{1}>L_{2}>L_{3}$, and the original plan choose using hash join. In order to obtain more bounty, the agent may choose to initially override the method at $O_{i}$ to a merge join, and then in subsequent steps, override it to a nested loop join. To address such issue, we calculate the minimum number of steps required from the original incomplete plan to the current incomplete plan, denoted as $minsteps(ICP_{t}^{e})$. Additionally, we introduce a penalty coefficient $\gamma>0$ to represent the weight of penalty in reward. We set the penalty value according to the following formula:
@@
. In the case of the initial step, it takes the query $Q$ as input and outputs a complete plan $CP$. In non-initial steps, it takes both the incomplete plan $ICP$ and the query $Q$ as input, generating a complete plan $CP$ steered by $ICP$. We then encode the $CP$ to obtain the new state by using (1). For ii), after obtaining $CP$, we provide a reward as described in the part of Reward. The key difference is that in the real environment, we use the DBMS executor $\Psi_{p}$ to execute the plan and then use the latency to represent the performance of plan $U(CP)$. Then we calculate the advantage and the final reward. In the simulated environment, the advantage value is evaluated by asymmetric advantage model without executing the plan. The specific details will be discussed in IV-B.
@@
2:     $CP^{e}_{0}(CP^{Q}_{ORI})\leftarrow\Gamma_{p}(Q,/)$
3:     $ICP^{e}_{0}\leftarrow Extract(CP^{e}_{0})$
4:     Episode Buffer Set $T=[ICP^{e}_{0}]$
5:     Estimated Optimal Plan $\overline{CP}^{e}_{0}\leftarrow CP^{e}_{0}$
@@
7:         $actionmask\leftarrow ICP^{e}_{t-1}$
8:         if $t>1\cap a^{e}_{t-1}\in[1,I_{s}]$ then
@@
11:         $s^{e}_{t-1}=State(CP^{e}_{t-1})$
12:         $statevec^{e}_{t-1}=\phi(s^{e}_{t-1})$
13:         $a^{e}_{t}=\pi(statevec^{e}_{t-1},actionmask)$
14:         $ICP^{e}_{t}=Act(a^{e}_{t},ICP^{e}_{t-1})$
15:         $CP^{e}_{t}\leftarrow\Gamma_{p}(Q,ICP^{e}_{t})$
16:         $r^{e}_{t}=Penalty^{e}_{t}$
17:         if $ICP^{e}_{t}\notin T$ then
18:              $r^{e}_{t}=r^{e}_{t}+Bounty^{e}_{t}$
19:              $Add\ ICP^{e}_{t}\ to\ T$
@@
21:         $\overline{CP}^{e}_{t}\leftarrow CP^{e}_{t}\text{\ \ if\ \ }Adv(\overline{CP}^{e}_{t-1},CP^{e}_{t})>0$
22:         $experiences\leftarrow Collect(s^{e}_{t-1},a^{e}_{t},r^{e}_{t},s^{e}_{t})$